OpenStudy (anonymous):

When a popular running shoe is priced at $70, The Shoe House will sell 15 pairs each week. Using a survey, they have determined that for each decrease of $2 in price, 3 additional pairs will be sold each week. What selling price will give a weekly revenue of $2250?

OpenStudy (anonymous):

\[\rm Revenue = Price \cdot Pairs~Sold\]\[2250 = (70 - 2 \cdot n) \cdot (15 + 3 \cdot n)\]We can solve for \(n\) such that the above equation is satisfied. Then use the following relation to determine the price of the shoes that generates the required revenue. \[{\rm Price } = 70 - 2 \cdot n\]

OpenStudy (anonymous):

so how do I solve for n?

OpenStudy (anonymous):

Foil, then use the quadratic equation.